OpenStudy (lilai3):

Brittany received 8 silver dollars on her eighth birthday. After her next birthday she had 15 and after the next she had 22. After her eleventh birthday she had 29 silver dollars. How many silver dollars will she have after her 16th birthday if her collection increases at the same rate every year?

OpenStudy (anonymous):

After her 16th birthday, she will have 64 silver dollars...

OpenStudy (lilai3):

okay can you explain or sumding????

OpenStudy (anonymous):

See, her silver medals are increasing at the rate of 7 per year... So, in 5 years, she will have 35 silver dollars.. Add 29 and 35 = 64 silver dollars...

OpenStudy (lilai3):

let's see what @lalaly hasta say :)

OpenStudy (lalaly):

i was typing the same thing @waterineyes said but u can do it like this she got 8 in her eighth birthday and 16-8=8 so 8 years after 8+7(8)=8+56=64

OpenStudy (lilai3):

okay thanks so this question is officially closed thank you to both @waterineyes and @lalaly !!!!!

OpenStudy (lalaly):

your welcome :D:D

OpenStudy (anonymous):

Welcome dear...

OpenStudy (lilai3):

qwoiuthkanksdkjf wierupoijdkfl;aksjdfthankyoupeopleverymcuhasdfl;kjas;dlkfjfdakl;kjf

OpenStudy (lalaly):

lol

OpenStudy (lilai3):

you see it right?